Circular No. 5987

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994N IN UGC 5695 M. Turatto, Padova, communicates his discovery of a new supernova

during observations on May 10.0 UT at La Silla with the European Southern Observatory 3.6-m telescope of SN 1993N in UGC 5695 (α = 10^{h}27^{m}.1, δ = +13^{o}16′, equinox 1950.0). The new object is located 1′′.8 east and 9′′.8 north of the galaxy’s nucleus and 8′′.5 east of SN 1993N. Both supernovae entered the spectrograph slits oriented east-west; EFOSC spectra, preliminarily reduced by C. Gouiffes (Saclay), show that SN 1994N is of type II, ∼ 1–2 months past maximum, with typical features of H, Ca II, Na I, Ba II, Sc II, Fe I, and Fe II. The expansion velocity deduced from the P-Cyg absorption of Hα is ∼ 4500 km/s. B. Leibundgut measured magnitude R = 17.5 ± 0.5 with the 0.9-m Dutch telescope at ESO under poor conditions.

COMET TAKAMIZAWA (1994i)

Improved parabolic orbital elements from MPEC 1994-J06:

T = 1994 June 28.877 TT  ω = 191.^{o}693
Ω = 51.116 }2000.0
q = 1.95214 AU  i = 136.002
1994TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
May 8 17^{h}00.^{m}30 − 4^{o}52.^{′}8 1.126 2.053 148.^{o}0 15.^{o}1 10.4
10 16 48.89 − 5 21.2 1.097 2.046 152.3 13.3 10.3
12 16 36.78 − 5 51.0 1.072 2.039 156.6 11.4 10.2
14 16 24.04 − 6 21.7 1.050 2.032 160.7  9.5 10.2
16 16 10.73 − 6 53.1 1.033 2.025 164.4  7.7 10.1
18 15 56.96 − 7 24.7 1.020 2.019 167.1  6.4 10.1
20 15 42.85 − 7 56.0 1.012 2.013 168.0  6.0 10.1
22 15 28.56 − 8 26.6 1.008 2.007 166.6  6.7 10.0
24 15 14.22 − 8 56.0 1.010 2.001 163.4  8.3 10.0
26 14 59.99 − 9 23.9 1.016 1.996 159.4 10.3 10.0
28 14 46.02 − 9 49.9 1.026 1.991 154.9 12.5 10.0

Total magnitude estimates: May 8.34 UT, 10.4 (C. E. Spratt, Victoria, BC, 0.10-m refractor; independent visual discovery); 9.90, 11.2 (H. Mikuž, Ljubljana, Slovenia, 0.20-m Baker-Schmidt camera + V filter + CCD).

1994 May 10 Daniel W. E. Green